# Probabilistic Kernel Function for Fast Angle Testing

### v2

@@ Abstract
In this paper, we study the angle testing problem in the context of similarity search in high-dimensional Euclidean spaces and propose two projection-based probabilistic kernel functions, one designed for angle comparison and the other for angle thresholding. Unlike existing approaches that rely on random projection vectors drawn from Gaussian distributions, our approach leverages reference angles and employs a deterministic structure for the projection vectors. Notably, our kernel functions do not require asymptotic assumptions, such as the number of projection vectors tending to infinity, and can be both theoretically and experimentally shown to outperform Gaussian-distribution-based kernel functions. We apply the proposed kernel function to Approximate Nearest Neighbor Search (ANNS) and demonstrate that our approach achieves a 2.5X $\sim$ 3X higher query-per-second (QPS) throughput compared to the widely-used graph-based search algorithm HNSW.
@@ 1 Introduction
Among the various similarity measures for high-dimensional vectors, the $\ell_{2}$ norm, cosine distance, and inner product are the most commonly used in practice. As discussed in Yan et al. (2018); Dai et al. (2020); Lu et al. (2024), it is often possible to pre-compute and store the norms of vectors in advance, allowing these measures to be reduced to the computation of the cosine of the angle between two normalized vectors, thereby highlighting the central role of angle computation. On the other hand, in many real-world scenarios, we are not concerned with the exact values of the angles but rather with the outcome—which one is greater—of an angle comparison, which is referred to as the angle testing: Given a query vector $\bm{q}$ and data vectors $\bm{v_{1}}$, $\bm{v_{2}}$, $\bm{v}$ on sphere $\mathbb{S}^{d-1}$, typical operations include comparing $\left\langle\bm{q},\bm{v_{1}}\right\rangle$ and $\left\langle\bm{q},\bm{v_{2}}\right\rangle$, or determining whether $\left\langle\bm{q},\bm{v}\right\rangle$ exceeds a certain threshold. These operations, however, require computing exact cosines of angles, which have a cost of $O(d)$ per comparison and become expensive in high dimensions. To address this, we aim to design a computation-efficient probabilistic kernel function $K$ that can approximate these comparisons with reduced cost and high success probability. More precisely, we focus on the following two problems:
@@
(Probabilistic kernel function for comparison) Design a probabilistic kernel function $K$: $\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}\rightarrow RV$ with computational cost $o(d)$, where $RV$ denotes the set of random variables, such that, for any data vectors $v_{1},v_{2}\in\mathbb{S}^{d-1}$ and query $q\in\mathbb{S}^{d-1}$ satisfying $\left\langle q,v_{1}\right\rangle>\left\langle q,v_{2}\right\rangle$, we have $\Pr[K(q,v_{1})>K(q,v_{2})]>1-\epsilon$, where $\epsilon\leq 0.5$.
@@
(Probabilistic kernel function for thresholding) Given a fixed angle threshold $\theta\in(0,\pi)$, design a probabilistic kernel function $K$: $\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}\rightarrow RV$ with computational cost $o(d)$ such that for any $q,v_{1},v_{2}\in\mathbb{S}^{d-1}$ with angles $\phi_{1}<\theta$ between $q$ and $v_{1}$, and $\phi_{2}>\theta$ between $q$ and $v_{2}$, we have $\Pr[K(q,v_{1})>\cos\theta]\geq 1-\epsilon_{1}$, and $\Pr[K(q,v_{2})>\cos\theta]<\epsilon_{2}$, where $\epsilon_{1},\epsilon_{2}\leq 0.5$.
One of the main goals of this paper is to design appropriate $K$’s to solve the above two problems. Before proceeding, we note that much work has focused on the estimation of $\left\langle\bm{q},\bm{v}\right\rangle$, such as Johnson-Lindenstrauss (JL) bound (Johnson et al., 1984) and quantization-based techniques (Jégou et al., 2011; Ge et al., 2014; Martinez et al., 2018; Gao & Long, 2024; Guo et al., 2020). In contrast, the defined problems focus on comparing $\left\langle\bm{q},\bm{v}\right\rangle$ with another inner product or a threshold, necessitating a distinct theoretical analysis centered on probabilistic decision-making. Beyond the theoretical perspective, these two problems also give rise to a range of practical applications. The goal of Problem 1.1 aligns with that of the random-projection-based technique CEOs (Pham, 2021) under cosine distance (we postpone the general case of inner product to Sec. 6), allowing the designed kernel function to be applied to tasks where CEOs is effective, such as Maximum Inner Product Search (MIPS) (Pham, 2021), filtering of NN candidates (Pham & Liu, 2022), DBSCAN (Xu & Pham, 2024), and more. Notably, the goal of Problem 1.2 is similar to that of the graph-based ANNS approach PEOs (Lu et al., 2024), making the corresponding kernel function well suited for probabilistic routing tests in similarity graphs, which have demonstrated significant performance improvements over baseline graphs such as HNSW (Malkov & Yashunin, 2020). In Sec. 6, we will further elaborate on the applications of these two probabilistic kernel functions.
Despite addressing different tasks, all the techniques (Pham, 2021; Pham & Liu, 2022; Xu & Pham, 2024; Lu et al., 2024) mentioned above use Gaussian distribution to generate projection vectors and are built upon a common statistical result, as follows.
@@
(Theorem 1 in Pham (2021)) Given two vectors $\bm{v}$, $\bm{q}$ on $\mathbb{S}^{d-1}$, and $m$ random vectors $\{\bm{u}_{i}\}^{m}_{i=1}\sim\mathcal{N}(0,I^{d})$, let $\bm{u}_{\mathrm{max}}={\mathop{\rm argmax}_{\bm{u_{i}}}}|\bm{q}^{\top}\bm{u_{i}}|$. As $m$ goes infinity, we have:
@@
Lemma 1.3 builds the relationship between angles and corresponding projection vectors. Actually, $\bm{v}^{\top}\bm{u_{\text{max}}}$ can be viewed as an indicator of the cosine angle $\bm{q}^{\top}\bm{v}$. More specifically, the larger $\bm{v}^{\top}\bm{u_{\text{max}}}$ is, the more likely it is that $\bm{q}^{\top}\bm{v}$ is large. On the other hand, $\bm{v}^{\top}\bm{u_{\text{max}}}$ can be computed beforehand during the indexing phase and can be easily accessed during the query phase, making $K(\bm{q},\bm{v})=\bm{v}^{\top}\bm{u_{\text{max}}}$ a suitable kernel function for various angle testing problems, e.g., Problem 1.1.
However, Lemma 1.3 has a significant theoretical limitation: Relationship (1) relies on the assumption that the number of projection vectors $m$ tends to infinity. Since the evaluation time of projection vectors depends on $m$, $m$ cannot be very large in practice. Moreover, since Pham & Liu (2022); Xu & Pham (2024); Lu et al. (2024) all used Lemma 1.3 to derive their theoretical results, these results are also affected by this limitation, and the impact of $m$ becomes even harder to predict in these applications.
@@
(3) Based on $K^{1}_{S}$, we propose a random-projection technique KS1 which can be used for CEOs-based tasks (Pham, 2021; Pham & Liu, 2022; Xu & Pham, 2024). Based on $K^{2}_{S}$, we introduce a new routing test called the KS2 test which can be used to accelerate the graph-based Approximate Nearest Neighbor Search (ANNS) (Lu et al., 2024) (Sec. 6).
(4) We experimentally show that KS1 provides a slight accuracy improvement (up to 0.8%) over CEOs. For ANNS, we show that HNSW+KS2 improves the query-per-second (QPS) of the state-of-the-art approach HNSW+PEOs (Lu et al., 2024) by 10% – 30%, along with a 5% reduction in index size.
@@ 2 Related Work
In high-dimensional Euclidean spaces, the estimation of angles via random-projection techniques, especially Locality Sensitive Hashing (LSH) (Indyk & Motwani, 1998; Andoni & Indyk, ; Andoni & Indyk, 2008), has a relatively long history. A classical LSH technique is SimHash (Charikar, 2002), whose basic idea is to generate multiple hyperplanes and partition the original space into many cells such that two vectors falling into the same cell are likely to have a small angle between them. Andoni et al. (2015) proposed a different LSH method called Falconn for angular distance, whose basic idea is to find the closest or furthest projection vector to the data vector and record this projection vector as a hash value, leading to better search performance than SimHash. Later, Pham (2021) employed Concomitants of Extreme Order Statistics (CEOs) to identify the projection with the largest or smallest inner product with the data vector, as shown in Lemma 1.3, and recorded the corresponding maximum or minimum projected value to obtain a more accurate estimation than using a hash value alone (Pham & Liu, 2022).
Due to its ease of implementation, CEOs has been employed in several similarity search tasks (Pham, 2021; Andoni et al., 2015; Xu & Pham, 2024), as mentioned in Sec. 1. Additionally, CEOs has been used to accelerate similarity graphs, which are among the leading structures for Approximate Nearest Neighbor Search (ANNS). By swapping the roles of query and data vectors in CEOs, Lu et al. (2024) introduced a space-partitioning technique and proposed the PEOs test, which can be used to compare the objective angle with a fixed threshold under probabilistic guarantees. This test was incorporated into the routing mechanisms of similarity graphs and achieved significant search performance improvements over original graph structures like HNSW (Malkov & Yashunin, 2020) and NSSG (Fu et al., 2022).
@@ 3 Two Probabilistic Kernel Functions
In Sec. 3, we aim to propose probabilistic kernel functions for Problems 1.1 and 1.2. First, we introduce some notation. Frequently used symbols in this paper are listed in Table 2 in Appendix. Let $\mathbb{R}^{d}$ be the ambient vector space. Define $H\in SO(d)\in\mathbb{R}^{d\times d}$ as a random rotation matrix [^1] Note that the definition here differs from that in Andoni et al. (2015), where the so-called random rotation matrix is actually a matrix with i.i.d. Gaussian entries) and let $S=[\bm{u_{1}},\bm{u_{2}},\ldots,\bm{u_{m}}]\in\mathbb{R}^{d\times m}$ be an arbitrary fixed set of $m$ points on the unit sphere $\mathbb{S}^{d-1}$. For any vector $\bm{v}\in\mathbb{S}^{d-1}$, define the reference vector of $\bm{v}$ with respect to $S$ as $Z_{S}(\bm{v})=\mathop{\rm argmax}_{\bm{u}\in S}\left\langle\bm{u},\bm{v}\right\rangle$. Let $A_{S}(\bm{v})$ denote the cosine of the reference angle with respect to $\bm{v}$, that is, $A_{S}(\bm{v})=\left\langle\bm{v},Z_{S}(\bm{v})\right\rangle$ . Next, we introduce two probabilistic kernel functions $K^{1}_{S}(\cdot,\cdot)$ and $K^{2}_{S}(\cdot,\cdot)$ as follows, where $K^{1}_{S}(\cdot,\cdot)$ corresponds to Problem 1.1 and $K^{2}_{S}(\cdot,\cdot)$ corresponds to Problem 1.2.
@@
Remarks. (1) (Exploitation of reference angle) In the design of existing projection techniques such as CEOs (Pham, 2021), Falconn (Andoni et al., 2015), Falconn++ (Pham & Liu, 2022), etc., only the reference vector $Z_{S}{(\cdot))}$ is utilized. In contrast, our kernel functions defined in Eq. (2) and Eq. (3) incorporate not only the reference vector $Z_{S}{(\cdot))}$ but also the reference angle information $A_{S}(\cdot)$ (although the reference angle is not explicitly shown in Eq. (2), its influence will become clear in Lemma 4.2). In fact, the reference angle plays a central role, as it is the key factor controlling the precision of angle estimation (see Lemma 4.2).
(2) (Generalizations of existing works) These two kernel functions can also be regarded as generalizations of CEOs and PEOs respectively in a certain sense. Specifically, if $S$ is taken as a point set generated via a Gaussian distribution and $Z_{S}(\bm{v})$ is replaced by the reference vector having the maximum inner product with the query, then $K^{1}_{S}(\bm{q},\bm{v})$ equals the indicator $\bm{v}^{\top}\bm{u_{\text{max}}}$ used in CEOs. Similarly, if we remove the term $A_{S}(H\bm{v})$ and take $S$ to be the same space-partitioned structure as that of PEOs, $K^{2}_{S}(\bm{q},\bm{v})$ is similar to the indicator of PEOs. In addition, Eq. (3) reduces to the estimator in Gao & Long (2024) when $S$ is taken to be the hypercube with scalar quantization.
(3) (Configuration of projection vectors) Although we do not currently require any specific properties of the configuration of $S$, it is clear that the configuration of $S$ impacts both $K^{1}_{S}(\bm{q},\bm{v})$ and $K^{2}_{S}(\bm{q},\bm{v})$. We will discuss the structure of $S$ in detail in Sec. 5. Notably, we will see that neither the hypercube adopted in Gao & Long (2024) nor the Gaussian distribution adopted in Pham (2021); Lu et al. (2024) provides the optimal configuration for projection vectors.

### v3

# arXiv:2505.20274v3[cs.LG] 01 Mar 2026

@@ Abstract
[…]
@@ 1 Introduction
Among the various similarity measures for high-dimensional vectors, the $\ell_{2}$ norm, cosine similarity, and inner product are the most commonly used in practice. As discussed in Yan et al. (2018); Dai et al. (2020); Lu et al. (2024), it is often possible to pre-compute and store the norms of vectors in advance, allowing these measures to be reduced to the computation of the cosine of the angle between two normalized vectors, thereby highlighting the central role of angle computation. On the other hand, in many real-world scenarios, we are not concerned with the exact values of the angles but rather with the outcome—which one is greater—of an angle comparison, which is referred to as angle testing: Given a query vector $\bm{q}$ and data vectors $\bm{v_{1}}$, $\bm{v_{2}}$, $\bm{v}$ on the sphere $\mathbb{S}^{d-1}$, typical operations include comparing $\left\langle\bm{q},\bm{v_{1}}\right\rangle$ and $\left\langle\bm{q},\bm{v_{2}}\right\rangle$, or determining whether $\left\langle\bm{q},\bm{v}\right\rangle$ exceeds a certain threshold. These operations, however, require computing exact cosines of angles, which has a cost of $O(d)$ per comparison and becomes expensive in high dimensions. To address this, we aim to design a computationally efficient probabilistic kernel function $K$ that can approximate these comparisons with reduced cost and high success probability. More precisely, we focus on the following two problems:
@@
(Probabilistic kernel function for comparison) Design a probabilistic kernel function $K$: $\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}\rightarrow RV$ with computational cost $o(d)$, where $RV$ denotes the set of random variables, such that, for any data vectors $\bm{v_{1}},\bm{v_{2}}\in\mathbb{S}^{d-1}$ and query $\bm{q}\in\mathbb{S}^{d-1}$ satisfying $\left\langle\bm{q},\bm{v_{1}}\right\rangle>\left\langle\bm{q},\bm{v_{2}}\right\rangle$, we have $\Pr[K(\bm{q},\bm{v_{1}})>K(\bm{q},\bm{v_{2}})]>1-\epsilon$, where $\epsilon\leq 0.5$.
@@
(Probabilistic kernel function for thresholding) Given a fixed angle threshold $\theta\in(0,\pi)$, design a probabilistic kernel function $K$: $\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}\rightarrow RV$ with computational cost $o(d)$ such that for any $\bm{q},\bm{v_{1}},\bm{v_{2}}\in\mathbb{S}^{d-1}$ with angles $\phi_{1}<\theta$ between $\bm{q}$ and $\bm{v_{1}}$, and $\phi_{2}>\theta$ between $\bm{q}$ and $\bm{v_{2}}$, we have $\Pr[K(\bm{q},\bm{v_{1}})>\cos\theta]\geq 1-\epsilon_{1}$, and $\Pr[K(\bm{q},\bm{v_{2}})>\cos\theta]<\epsilon_{2}$, where $\epsilon_{1},\epsilon_{2}\leq 0.5$.
One of the main goals of this paper is to design appropriate $K$’s to solve the above two problems. Before proceeding, we note that much work has focused on the estimation of $\left\langle\bm{q},\bm{v}\right\rangle$, such as Johnson-Lindenstrauss (JL) bound (Johnson et al., 1984) and quantization-based techniques (Jégou et al., 2011; Ge et al., 2014; Martinez et al., 2018; Gao and Long, 2024; Guo et al., 2020). In contrast, the defined problems focus on comparing $\left\langle\bm{q},\bm{v}\right\rangle$ with another inner product or a threshold, necessitating a distinct theoretical analysis centered on probabilistic decision-making. Beyond the theoretical perspective, these two problems also give rise to a range of practical applications. The goal of Problem 1.1 aligns with that of the random-projection-based technique CEOs (Pham, 2021) under cosine similarity (we postpone the general case of inner product to Sec. 6), allowing the designed kernel function to be applied to tasks where CEOs is effective, such as Maximum Inner Product Search (MIPS) (Pham, 2021), filtering of NN candidates (Pham and Liu, 2022), DBSCAN (Xu and Pham, 2024), and more. Notably, the goal of Problem 1.2 is similar to that of the graph-based ANNS approach PEOs (Lu et al., 2024), making the corresponding kernel function well suited for probabilistic routing tests in similarity graphs, which have demonstrated significant performance improvements over baseline graphs such as HNSW (Malkov and Yashunin, 2020). In Sec. 6, we will further elaborate on the applications of these two probabilistic kernel functions.
Despite addressing different tasks, all the techniques (Pham, 2021; Pham and Liu, 2022; Xu and Pham, 2024; Lu et al., 2024) mentioned above use Gaussian distribution to generate projection vectors and are built upon a common statistical result, as follows.
@@
(Theorem 1 in Pham (2021)) Given two vectors $\bm{v}$, $\bm{q}$ on $\mathbb{S}^{d-1}$, and $m$ random vectors $\{\bm{u}_{i}\}^{m}_{i=1}\sim\mathcal{N}(0,I^{d})$, let $\bm{u}_{\mathrm{max}}={\mathop{\rm argmax}_{\bm{u_{i}}}}|\bm{q}^{\top}\bm{u_{i}}|$. As $m$ goes to infinity, we have:
@@
Lemma 1.3 builds the relationship between angles and corresponding projection vectors. Actually, $\bm{v}^{\top}\bm{u_{\text{max}}}$ can be viewed as an indicator of the cosine of the angle between $\bm{q}$ and $\bm{v}$. More specifically, the larger $\bm{v}^{\top}\bm{u_{\text{max}}}$ is, the more likely it is that $\bm{q}^{\top}\bm{v}$ is large. On the other hand, all $\bm{v}^{\top}\bm{u_{i}}$’s can be computed beforehand during the indexing phase and $\bm{v}^{\top}\bm{u_{\max}}$ can be easily accessed during the query phase, making $K(\bm{q},\bm{v})=\bm{v}^{\top}\bm{u_{\text{max}}}$ a suitable kernel function for various angle testing problems, e.g., Problem 1.1.
However, Lemma 1.3 has a significant theoretical limitation: Relationship (1) relies on the assumption that the number of projection vectors $m$ tends to infinity. Since the computational cost of evaluating projection vectors grows with $m$, $m$ cannot be very large in practice. Moreover, since Pham and Liu (2022); Xu and Pham (2024); Lu et al. (2024) all used Lemma 1.3 to derive their theoretical results, these results are also affected by this limitation, and the impact of $m$ becomes even harder to predict in these applications.
@@
(3) Based on $K^{1}_{S}$, we propose a random-projection technique KS1 which can be used for CEOs-based tasks (Pham, 2021; Pham and Liu, 2022; Xu and Pham, 2024). Based on $K^{2}_{S}$, we introduce a new routing test called the KS2 test which can be used to accelerate the graph-based Approximate Nearest Neighbor Search (ANNS) (Lu et al., 2024) (Sec. 6).
[…]
@@ 2 Related Work
In high-dimensional Euclidean spaces, the estimation of angles via random-projection techniques, especially Locality Sensitive Hashing (LSH) (Indyk and Motwani, 1998; Andoni and Indyk, 2005; 2008), has a relatively long history. A classical LSH technique is SimHash (Charikar, 2002), whose basic idea is to generate multiple hyperplanes and partition the original space into many cells such that two vectors falling into the same cell are likely to have a small angle between them. Andoni et al. (2015) proposed a different LSH method called Falconn for angular distance, whose basic idea is to find the closest or furthest projection vector to the data vector and record this projection vector as a hash value, leading to better search performance than SimHash. Later, Pham (2021) employed Concomitants of Extreme Order Statistics (CEOs) to identify the projection with the largest or smallest inner product with the data vector, as shown in Lemma 1.3, and recorded the corresponding maximum or minimum projected value to obtain a more accurate estimation than using a hash value alone (Pham and Liu, 2022).
Due to its ease of implementation, CEOs has been employed in several similarity search tasks (Pham, 2021; Andoni et al., 2015; Xu and Pham, 2024), as mentioned in Sec. 1. Additionally, CEOs has been used to accelerate similarity graphs, which are among the leading structures for Approximate Nearest Neighbor Search (ANNS). By swapping the roles of query and data vectors in CEOs, Lu et al. (2024) introduced a space-partitioning technique and proposed the PEOs test, which can be used to compare the objective angle with a fixed threshold under probabilistic guarantees. This test was incorporated into the routing mechanisms of similarity graphs and achieved significant search performance improvements over original graph structures like HNSW (Malkov and Yashunin, 2020) and NSSG (Fu et al., 2022).
@@ 3 Two Probabilistic Kernel Functions
In Sec. 3, we aim to propose probabilistic kernel functions for Problems 1.1 and 1.2. First, we introduce some notation. Frequently used symbols in this paper are listed in Table 2 in Appendix. Let $\mathbb{R}^{d}$ be the ambient vector space. Define $H\in SO(d)\subset\mathbb{R}^{d\times d}$ as a random rotation matrix, where $SO(d)$ denotes the special orthogonal group of $d\times d$ rotation matrices [^1] Note that the definition here differs from that in Andoni et al. (2015), where the so-called random rotation matrix is actually a matrix with i.i.d. Gaussian entries.. Let $S=[\bm{u_{1}},\bm{u_{2}},\ldots,\bm{u_{m}}]\in\mathbb{R}^{d\times m}$ be an arbitrary fixed set of $m$ points on the unit sphere $\mathbb{S}^{d-1}$. For any vector $\bm{v}\in\mathbb{S}^{d-1}$, define the reference vector of $\bm{v}$ with respect to $S$ as $Z_{S}(\bm{v})=\mathop{\rm argmax}_{\bm{u}\in S}\left\langle\bm{u},\bm{v}\right\rangle$. Let $A_{S}(\bm{v})$ denote the inner product with the reference vector with respect to $\bm{v}$, that is, $A_{S}(\bm{v})=\left\langle\bm{v},Z_{S}(\bm{v})\right\rangle$. Let $HS=\{Hu:u\in S\}$ denote the rotated configuration of $S$. Next, we introduce two probabilistic kernel functions $K^{1}_{S}(\cdot,\cdot)$ and $K^{2}_{S}(\cdot,\cdot)$ as follows, where $K^{1}_{S}(\cdot,\cdot)$ corresponds to Problem 1.1 and $K^{2}_{S}(\cdot,\cdot)$ corresponds to Problem 1.2.
@@
Remarks. (1) (Exploitation of reference angle) In the design of existing projection techniques such as CEOs (Pham, 2021), Falconn (Andoni et al., 2015), Falconn++ (Pham and Liu, 2022), etc., only the reference vector $Z_{S}{(\cdot)}$ is utilized. In contrast, our kernel functions defined in Eq. (2) and Eq. (3) incorporate not only the reference vector $Z_{S}{(\cdot)}$ but also the reference angle information $A_{S}(\cdot)$ (although the reference angle is not explicitly shown in Eq. (2), its influence will become clear in Lemma 4.2). In fact, the reference angle plays a central role, as it is the key factor controlling the precision of angle estimation (see Lemma 4.2).
(2) (Generalizations of existing works) These two kernel functions can also be regarded as generalizations of CEOs and PEOs respectively in a certain sense. Specifically, if $S$ is taken as a point set generated via a Gaussian distribution and $Z_{S}(\bm{v})$ is replaced by the reference vector having the maximum inner product with the query, then $K^{1}_{S}(\bm{q},\bm{v})$ equals the indicator $\bm{v}^{\top}\bm{u_{\text{max}}}$ used in CEOs. Similarly, if we remove the term $A_{S}(H\bm{v})$ and take $S$ to be the same space-partitioned structure as that of PEOs, $K^{2}_{S}(\bm{q},\bm{v})$ is similar to the indicator of PEOs. In addition, Eq. (3) reduces to the estimator in Gao and Long (2024) when $S$ is taken to be the hypercube with scalar quantization.
(3) (Configuration of projection vectors) Although we do not currently require any specific properties of the configuration of $S$, it is clear that the configuration of $S$ impacts both $K^{1}_{S}(\bm{q},\bm{v})$ and $K^{2}_{S}(\bm{q},\bm{v})$. We will discuss the structure of $S$ in detail in Sec. 5. Notably, we will see that neither the hypercube adopted in Gao and Long (2024) nor the Gaussian distribution adopted in Pham (2021); Lu et al. (2024) provides the optimal configuration for projection vectors.